(a) Verify that y = e2xsinx is a solution to the differential equation d2y/dx2 - 4dy/dx+5y=0

(b) Differentiate the following functions with respect to x: (i) ln(1 + sin2 x) (ii) xx.
(c) Evaluate the integral 2x3 −4x−8 /x4 −x3 +4x2 −4x dx.
1
Expert's answer
2020-06-28T18:28:25-0400

(a) "y(x)=e^{2x} sin x."

Let's differentiate this function.

"\frac{dy}{dx}=(e^{2x})^\prime sin x+e^{2x}(sin x)^\prime=2e^{2x}sin x+e^{2x}cos x."

"\frac{d^2y}{dx^2}=2(e^{2x})^\prime sin x+2e^{2x}(sin x)^\prime+(e^{2x})^\prime cos x+ e^{2x}(cos x)^\prime="

"=4e^{2x} sin x+2e^{2x}cos x+ 2e^{2x}cos x-e^{2x} sin x=3e^{2x} sin x+4e^{2x}cos x."

Hence we have:

"\frac{d^2y}{dx^2} - 4\frac{dy}{dx}+5y=3e^{2x} sin x+4e^{2x}cos x-4(2e^{2x}sin x+e^{2x}cos x)+5e^{2x}sin x="

"=3e^{2x} sin x+4e^{2x}cos x-8e^{2x}sin x-4e^{2x}cos x+5e^{2x}sin x=0,"

so that the function "y(x)=e^{2x} sin x" is a solution to the differential equation "\frac{d^2y}{dx^2} - 4\frac{dy}{dx}+5y=0."

(b)

  • (i) "(ln(1 + sin^2x))^\prime =\frac{1}{1+sin^2x}(1+sin^2x)^\prime="

"\frac{1}{1+sin^2x}2sin xcos x=\frac{sin 2x}{1+sin^2x}."

  • (ii) "(x^x)^\prime=(e^{x ln x})^\prime=e^{x ln x}(x ln x)^\prime=e^{x ln x}((x)^\prime ln x+ x (ln x)^\prime)="

"=e^{x ln x}(1ln x+x\frac{1}{x})=e^{x ln x}(ln x+1)=x^x (lnx+1)."

(c) We'll use the method of partial fractions to evaluate the integral "\int \frac{2x^3 −4x−8}{x^4 −x^3 +4x^2 −4x} dx."

"\frac{2x^3 −4x−8}{x^4 −x^3 +4x^2 −4x}=\frac{2x^3 −4x−8}{x^3(x-1)+4x(x-1)}= \frac{2x^3 −4x−8}{(x^3+4x)(x-1)}="

"=\frac{2x^3 −4x−8}{x(x-1)(x^2+4)}=\frac{A}{x}+\frac{B}{x-1}+\frac{Cx+D}{x^2+4}."

Then we have:

"A(x-1)(x^2+4)+Bx(x^2+4)+(Cx+D)x(x-1)=2x^3-4x-8."

"A(x^3+4x-x^2-4)+B(x^3+4x)+C(x^3-x^2)+D(x^2-x)=2x^3-4x-8."

"(A+B+C)x^3+(-A-C+D)x^2+(4A+4B-D)x+(-4A)=2x^3-4x-8."

Hence:

"A+B+C=2,"

"\u2212A\u2212C+D=0,"

"4A+4B\u2212D=\u22124,"

"\u22124A=\u22128."

From the last equation: "A=2". We have:

"B+C=0, D-C=2, 4B-D=-12."

Then "B=-C, D=C+2,"

and we have the next equation: "-4C-2-C=-12,"

"-5C=-10, C=2."

Then "B=-C=-2, D=C+2=2+2=4."

So we have:

"\frac{2x^3 −4x−8}{x^4 −x^3 +4x^2 −4x}=\frac{2}{x}-\frac{2}{x-1}+\frac{2x+4}{x^2+4}."

Now we can evaluate the integral:

"\int \frac{2x^3 −4x−8}{x^4 −x^3 +4x^2 −4x} dx=\int(\frac{2}{x}-\frac{2}{x-1}+\frac{2x+4}{x^2+4})dx="

"=\int\frac{2}{x}dx-\int\frac{2}{x-1}dx+\int\frac{2x+4}{x^2+4}dx=\int\frac{2}{x}dx-\int\frac{2}{x-1}dx+\int\frac{2x}{x^2+4}dx+\int\frac{4}{x^2+4}dx="

"=\int\frac{2}{x}dx-\int\frac{2}{x-1}dx+\int\frac{d(x^2+4)}{x^2+4}+\int\frac{4}{x^2+4}dx="

"=2ln|x|-2ln|x-1|+ln|x^2+4|+4\cdot \frac{1}{2}arctan\frac{x}{2}+C="

"=2ln|x|-2ln|x-1|+ln|x^2+4|+2arctan\frac{x}{2}+C."